he Caraway Seed Company grows heirloom tomatoes and sells their seeds. The heirloom tomato plants are preferred by many growers for their superior flavor. At the end of the most recent year the firm had current assets of $ 49,200 net fixed assets of $ 252,000, current liabilities of $ 30,200, and long-term debt of $99,500.

a. Calculate Caraway's stockholders' equity.

b. What is the firm's net working capital?

c. If Caraway's current liabilities consist of $20,200 in accounts payable and $10,000 in short-term debt (notes payable), what is the firm's net working capital?
Answer #1

Given that, The firm's

Net fixed assets = $252,000

Current Assets = $49,200

Current Liabilities = $30,200

Long term liabilities = $99,500

A) We know that,

Total Assets = Total Liabilities + Equity

Current Assets + Fixed assets = Current liabilities + Long term liabilities + Equity

49200 + 252000 = 30200 + 99500 + Equity

Equity = $171,500

B) Net working capital = Current assets - Current Liabilities

= 49200 - 30200

= $ 19200

C) Both Accounts payable and Short term debt will be paid in the current accounting cycle, hence will be a part of Current Liabilities.

Net working capital = Current assets - Current Liabilities

= 49200 - 30200

= $ 19200